find the slope of every line that is parallel to the graph of the equation x-8y=-1
Answer by fcabanski(1391)

Parallel lines have the same slope. To find the slope, write the equation in slope-intercept form y=mx+b where m=slope.


x-8y=-1 | Subtract x from both sides

-8y = -x - 1 | Divide both sides by -8

y = 1/8x + 1/8


The slope is 1/8, so the slope of all lines parallel to this line is 1/8.


Contact me for private tutoring - fcabanski@hotmail.com